OpenStudy (anonymous):

553+35=? I need COMPENSATION, please!

OpenStudy (anonymous):

588

OpenStudy (anonymous):

553 + 35 is (553 -3) + (35 +3) = 550 + 38 =588

OpenStudy (anonymous):

Well, I need compensation as if you -3 oof one then you have to =3 to the other!